OpenStudy (anonymous):

complete the square: x^2+4x-6=0

OpenStudy (anonymous):

i got √(x+2)^2= √10

OpenStudy (anonymous):

(x+2)^2-6-4=0 (x+2)^2=10 x=-2+\[\sqrt{?}\]10 x=-2-\[\sqrt{?}\]10

OpenStudy (anonymous):

thanks guys